OpenStudy (anonymous):

how do you find the y intercept in two ordered pairs. Ex. (2,40 and (4, 9) Not a real problem, but i need to know how to do it.

OpenStudy (anonymous):

Start by finding the slope. Do you know the slope formula?

OpenStudy (anonymous):

yeah, i found the slope, in my problem its 1/2

OpenStudy (anonymous):

The slope between (2,40) and (4,9) isn't 1/2. I assume you're working with two different points?

OpenStudy (anonymous):

no.. haha i was just wondering how to do this. My two ordered pairs in my problem are (2,120) and 4, 240)

OpenStudy (anonymous):

Ah, ok. So the next step is to take the slope you found, and one point and plug them both into the point slope form of a line.

OpenStudy (anonymous):

\[y-k = m(x-h)\] Where m is the slope, and (h,k) is one of the points.

OpenStudy (anonymous):

to find the y-intercept?

OpenStudy (anonymous):

yes.

OpenStudy (anonymous):

oh, thanks

OpenStudy (anonymous):

so what would the y intercept be? 6?

OpenStudy (anonymous):

Once you have the equation for the line you can find the y intercept by plugging in a 0 for x and solve for y. Or you can find the x intercept by plugging in a 0 for y and solving for x.

OpenStudy (anonymous):

\(y-120 = \frac{1}{2}(x - 2)\)\[\implies y = \frac{1}{2}x - 1 + 120\] Plugging in 0 for x.. \[\implies y = 0 -1 + 120 = 119\]

OpenStudy (anonymous):

Does that make sense?

OpenStudy (anonymous):

yes. alot actually. Thank you. :)